# On degenerate metric resolvent: local monotonicity, restricted maximality and beyond

### v2

@@ 2.2.2 $\mathcal{Q}$-firm nonexpansiveness of $\mathcal{T}$
Proof 2.7.
Let $\{x^{k}\}_{k\in\mathbb{N}}$ be a sequence in $D\subseteq\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{S}$, such that $x^{k}\rightharpoonup x$ and $x^{k}-\mathcal{S}x^{k}\rightarrow u$. Since $D$ is weakly sequentially closed, and $\mathsf{ran}\mathcal{Q}$ is closed (by Assumption 1), we have $x\in D$ and $u\in\mathsf{ran}\mathcal{Q}$. For every $k\in\mathbb{N}$, it follows from the $\mathcal{Q}$-nonexpansiveness of $\mathcal{S}$ that
$\big{\|}x-\mathcal{S}x-u\big{\|}_{\mathcal{Q}}^{2}$ (9)
$=$ $\big{\|}x^{k}-\mathcal{S}x-u\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x^{k}-x\big{\|}_{\mathcal{Q}}^{2}-2\big{\langle}x^{k}-x\big{|}x-\mathcal{S}x-u\big{\rangle}_{\mathcal{Q}}$
$=$ $\big{\|}x^{k}-\mathcal{S}x^{k}-u\big{\|}_{\mathcal{Q}}^{2}+2\big{\langle}x^{k}-\mathcal{S}x^{k}-u\big{|}\mathcal{S}x^{k}-\mathcal{S}x\big{\rangle}_{\mathcal{Q}}$
$+$ $\big{\|}\mathcal{S}x^{k}-\mathcal{S}x\big{\|}_{\mathcal{Q}}^{2}-\big{\|}x^{k}-x\big{\|}_{\mathcal{Q}}^{2}-2\big{\langle}x^{k}-x\big{|}x-\mathcal{S}x-u\big{\rangle}_{\mathcal{Q}}$
$\leq$ $\big{\|}x^{k}-\mathcal{S}x^{k}-u\big{\|}_{\mathcal{Q}}^{2}+2\big{\langle}x^{k}-\mathcal{S}x^{k}-u\big{|}\mathcal{S}x^{k}-\mathcal{S}x\big{\rangle}_{\mathcal{Q}}$
-- $2\big{\langle}x^{k}-x\big{|}x-\mathcal{S}x-u\big{\rangle}_{\mathcal{Q}}.$
By assumption, $x^{k}-\mathcal{S}x^{k}-u\rightarrow 0$ and $x^{k}-x\rightharpoonup 0$ as $k\rightarrow\infty$, and hence
$\mathcal{S}x^{k}-\mathcal{S}x=(\mathcal{S}x^{k}-x^{k})+(x^{k}-x)+x-\mathcal{S}x\rightharpoonup x-\mathcal{S}x-u.$
Taking the limit as $k\rightarrow\infty$ in (9) and appealing to Fact 2, we obtain $\|x-\mathcal{S}x-u\|_{\mathcal{Q}}=0$, which leads to $x-\mathcal{S}x=u$ since $x,\mathcal{S}x,u\in\mathsf{ran}\mathcal{Q}$.

### v3

@@ 2.2 Basic properties
[…]
Given $\mathcal{T}$ as (2) satisfying Assumptions 1 and 2, let $D$ be a nonempty closed subset of $\mathsf{ran}\mathcal{Q}\cap\mathsf{dom}\mathcal{T}$, then $\mathcal{I}-\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-demiclosed.
Proof 2.9.
In view of Lemma 2.7, it suffices to show that $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\circ\mathcal{T}|_{\mathsf{ran}\mathcal{Q}}:D\mapsto\mathsf{ran}\mathcal{Q}$ is $\mathcal{Q}$-nonexpansive. Indeed, this immediately follows from Fact 2.
2.3 Generalized Moreau’s decomposition identity
Lemma 2.10 develops some equivalent forms of $\mathcal{T}$, which generalize the classic Moreau’s decomposition identity [4, Propositioin 23.20], and will be used for proving Theorem 3.1. The proof is postponed to Appendix A.
Lemma 2.10.
Under Assumptions 1 and 2, the following equalities hold:
$\mathcal{Q}-\mathcal{Q}(\mathcal{A}+\mathcal{Q})^{-1}\mathcal{Q}$
[…]
where $\sqrt{\mathcal{Q}}$ is defined as Fact 1-(iii).
3 Full domain
3.1 Main result
Theorem 3.1 provides a sufficient and necessary condition on $\mathcal{A}$ and $\mathcal{Q}$ for the full domain of $\mathcal{T}$.
Theorem 3.1.
Under Assumption 1, $\mathcal{T}$ given as (2) has full domain, if and only if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ (or equivalently saying, $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$) $:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$ is maximally monotone.
Let us first discuss the maximal monotonicity of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$, from which then readily follows the proof of Theorem 3.1. This concept is essentially a restricted version of maximal monotonicity from $\mathcal{H}$ to $2^{\mathcal{H}}$, by confining both domain and range of the operator to $\mathsf{ran}\mathcal{Q}$. Indeed, this restricted maximal monotonicity is closely related to Definition 2.1, due to the following lemma.
Lemma 3.2.
The following 3 statements are equivalent.
(ii) $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone;
Proof 3.3.
(i)$\Longleftrightarrow$(ii): Proposition 2.2.
(ii)$\Longrightarrow$(iii): If $\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
(iii)$\Longrightarrow$(ii): If $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is monotone, then we have:
It is natural to develop a restricted version of classic Minty’s theorem [4, Theorem 21.1] for the restricted maximality of $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$.
Theorem 3.4 (Restricted Minty’s theorem).
The proof is exactly the same as [4, Theorem 21.1], if the subspace $\mathsf{ran}\mathcal{Q}$ is regarded as the ambient space. Actually, our specific $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ in Theorem 3.4 can be replaced by any monotone operator $\mathcal{B}:\mathsf{ran}\mathcal{Q}\mapsto 2^{\mathsf{ran}\mathcal{Q}}$, of course, including $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$.
We are now ready to prove Theorem 3.1.
Proof 3.5.
Combining Lemma 2.10-(4) and Theorem 3.4, we claim that $\mathcal{T}$ has full domain, if and only if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone. What remains to show is the equivalence of maximal monotonicity between $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ and $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$.
First, if $\mathcal{P}_{\mathsf{ran}\mathcal{Q}}\mathcal{A}^{-1}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have:
Conversely, if $\sqrt{\mathcal{Q}}\mathcal{A}^{-1}\sqrt{\mathcal{Q}}|_{\mathsf{ran}\mathcal{Q}}$ is maximally monotone, then for any given $(u,x)\in\mathsf{ran}\mathcal{Q}\times\mathsf{ran}\mathcal{Q}$, we have: